6- At the beginning of an effective red, vehicles are arriving at an approach at the rate of 600 veh/h and 18 vehicles are left in the queue from the previous cycle (at the end of the previous cycle’s effective green). However, due to the end of a major sporting event, the arrival rate is continuously increasing at a constant rate of 150 veh/h/min (after 1 minute the arrival rate will be 750 veh/h, after 2 minutes 900 veh/h, etc.). The saturation flow rate of the approach is 1800 veh/h, the cycle length is 60 seconds, and the effective green time is 40 seconds. Determine the total vehicle delay until complete queue clearance. (Assume D/D/1 queuing.)

can someone please help
Answer #1

Solution . Ciseen time = go seands Red time = to Beconde 100 Seconds Cycle length Amival Rate = Soo Vehler (0.138 Vehl bec) nProportion of vehicles Estopped. Ps xlot to) +18+9) 84.03 = 0.128 ( 70+ 14,01) 0.138 (70 +30) = 0.8403 Maximum number of vehiMaximum Delay of my whicle dmax r = Tos They Solution combe determined; with a = 0.84 Pooperties the cycle queue mumber of ve